# PGS: Pose-Guided Supervision for Mitigating Clothes-Changing in Person Re-Identification

### 31 Dec 2023 v2

@@ Abstract
[…]
Index Terms— Person Re-Identification, Pose Guidance, Clothes-Changing
@@ 1 Introduction
Person Re-Identification (Re-ID) is the task of tracking and distinguishing individuals across multiple cameras and viewpoints. The application of this technique is in surveillance camera systems, enabling the search, tracking, and identification of unidentified individuals within a town or specific area, thereby enhancing security measures.
Although deep learning methods have obtained amazing research progress in Re-ID, the clothes-changing problem is still a challenge that should be considered. This problem is especially critical for real-world applications where people always change their clothes. In recent years, researchers in the Re-ID field have increasingly focused on addressing the clothing-change problem, exemplified by the introduction of VC-Clothes [1] and LTCC [2]. The primary difficulty in tackling this challenge arises from deep learning models relying heavily on clothing appearance to make predictions rather than considering biometric traits (e.g., face, hair, pose, gait). Our analysis underscores that the absence of pose information during the learning process is a significant factor contributing to the clothing-change problem. Therefore, a promising strategy involves prioritizing pose knowledge consisting of information about different human parts, which can be extracted using frozen pre-trained pose estimation models.
As a result, we propose the Pose-Guided Supervision (PGS), a simple and innovative framework for the clothes-changing Re-ID task. In essence, our method builds upon three main modules: human encoder, pose encoder, and Pose-to-Human Projection module (PHP). The human encoder efficiently adapts SOLIDER [3], a self-supervised learning framework for human-centric tasks, to extract the general human representations from the input images. At the same time, the pose encoder is taken from the pose estimation model OpenPose [4] to generate meaningful global features for biometric identification, which includes the pose, body, foot, hand, and facial. Our PHP, which includes several projectors, operates between two encoders to transfer the pose information extracted from the pose encoder to the human encoder across various scales. In practice, we fine-tune the pre-trained human encoder and train our projectors while freezing the pose encoder since we only need the general human pose information from the pose encoder. In summary, this paper’s main contributions are as follows:
@@
We introduce a Re-ID framework called PGS, incorporating a frozen pre-trained pose estimation model to guide the human representation model across different scales.
We conduct comprehensive ablation studies on multiple datasets to assess the effectiveness of our proposed PGS framework.
@@ 2 Related Work
Person Re-Identification: Bag of Tricks method [5] serves as the baseline approach for subsequent studies since it provides effective training tricks on this task. ABDNet [6] integrates two attention modules (e.g., CAM and PAM) with the novel regularization to learn robust feature representation. TransReID [7] proposes a pure transformer-based network with two novel modules to mitigate the data bias toward various cameras or viewpoints. Auto-ReID+ [8] offers the multi-branch macro network search architecture that combines multi-scale features, local, and global features to benefit the ReID model. PGFL-KD [9] exploits the pose-guided interaction learning to regularize the global feature’s learning. AML [10] designs an adaptive multiple loss baseline on the base of triplet loss to avoid the misjudgment of samples.
@@
Clothes-changing Re-Identification: The introduction of the LTCC dataset [2] marked a significant development to address the clothes-changing problem. Subsequent research has produced various methods to tackle this challenge, FSAM [11] aims to get the coarse ID masks with structure-related details, incorporating ID-relevant information for discriminative structural feature extraction. CAL [12] proposes a novel multi-positive-class classification loss to formulate multi-class adversarial learning. GI-ReID [13] integrates the clothes classification and the casual inference model for mitigating the bias on the clothes information. AIM [14] defines a causality-based auto intervention model to mitigate clothing bias for the robust cloth-changing person. FIR [15] introduces the fine-grained feature mining and fine-grained attribute recomposition module to boost the learning of robust features. Although previous works achieved impressive results on the clothes-changing Re-ID, it is worth noting that these methods lack focusing on the crucial parts of the body, which is the salient information to distinguish a specific person.
Based on the preceding analysis, we have designed PGS to prioritize body part information. This allows the model to efficiently leverage data from various body regions, such as the head, arms, legs, and upper body, for individual identification. This approach helps reduce the model’s dependence on clothing-related information. Moreover, integrating the frozen pre-trained pose estimation model also avoids the increase of the computation cost in the inference stage.
3 Proposed PGS
In this section, we introduce our PGS framework, which comprises three main modules: the human encoder, the pose encoder, and the pose-to-human projection module (PHP). Unlike PGFL-KD [9], which involves designing complex modules for body part feature extraction, our method adopts a more straightforward strategy by leveraging frozen pre-trained features from a pose estimation model, as illustrated in Figure 1.
@@ 3 Proposed PGS
The human encoder takes in the image, learns the feature representations and unique human appearance information of each person, and then returns the image embedding. Meanwhile, the pose encoder (i.e., from a frozen pre-trained pose estimation model) is used to extract the human pose information from the input image and packs this information into a pose heatmap. In the training phase, the PHP encourages the human encoder to attain the human representation with additional pose information through a knowledge distillation mechanism. When inferring, only feature representations from the human encoder are used to accomplish the re-identification task. In other words, we leverage a frozen encoder pre-trained for pose estimation as a teacher to compel the student human encoder to retain pose information into the image embedding. We detail each component in the following subsections to provide more in-depth information.
@@ 3.3 Pose-to-Human Projection Module (PHP)
It is important to note that the human encoder outputs feature informative and semantically richer vectors than the pose encoder, which only contain the human pose information. From this insight, we can see that pose features cannot directly guide the features from the human encoder because if we do that, the human features will lose other vital information. Therefore, we design an indirect transfer knowledge mechanism (i.e., projector), which encourages the human encoder to learn pose information without the loss of any other information. Hence, intuitively, the projector serves as a non-linear, lightweight extractor, revealing human information within the higher-dimensional image representation of the human encoder and mapping it to the lower-dimensional pose heatmap space of the pose encoder. Subsequently, the resulting features at each stage are integrated into the guide loss for alignment with the pose feature obtained from the pose encoder.
@@ 3.4 Objective Functions
where is the triplet loss [16] to learn person-specific representations, and is the guide loss to transfer the pose knowledge from pose encoder to human encoder. The hyperparameter , which we set to 0.8, is used to control the contribution and effect of the guide loss on the whole system. See section 4.3 for the ablation study on the values.
@@ 4.1 Experimental Setup
Datasets: We evaluate the proposed PGS on five datasets. Market-1501 [19] contains 1,501 identities observed from 6 camera viewpoints, including 12,936 training images of 751 identities, 19,732 gallery images, and 2,228 queries. Duke-MTMC [20] comprises 34,183 images of 1,404 identities from eight cameras. It contains 16,522 training images, 17,661 gallery images, and 2,228 queries. CuHK03 [21] contains 14,097 pictures of 1,467 identities. Additionally, the dataset offers 20 different sets for training and testing, where 100 identities are reserved for testing in each set while the remaining identities are used for training. LTCC [2] contains 17,119 person images of 152 identities. The training set in this data includes 9,576 images with 77 identities (46 clothes-changing IDs and 31 clothes-consistent IDs). The testing set includes 493 query images and 7,050 test images with 75 identities (45 clothes-changing IDs + 30 clothes-consistent IDs). This dataset includes two types of testing: Standard (test cases including clothes-changing and clothes-consistent), and CC (test cases include only clothes-changing IDs). VC-Clothes [1] comprises synthesized human images. It encompasses a total of 19,060 images of 512 unique identities in four scenes. In the training phase, 9449 images of 256 identities are employed, leaving the remaining 9611 images of 256 identities for the testing phase.
@@
Implementation Detail: During the training and testing phase, all of the images are resized to dimensions of . Our training and testing for the clothes-consistent follow the setting of Bag of Trick [5], while in the clothes-changing testing, we follow the setup of LTCC [2]. The Pytorch framework is used, and all experiments are done on the Tesla A100 40GB GPU. For the hyperparameters, the batch size for the training and testing is 64; we also use the learning rate scheduler to support our training experiments with a base learning rate of 8e-4, and the temperature value equals 2. We choose the AdamW[22] as our optimizer, and the best weight has been gotten after 250 epochs. For data augmentation, we employ random horizontal flip and random erasing.
@@
Baselines: Based on two kinds of Re-ID problems, we select different baselines in two scenarios as follows. For clothes-consistent setting, we evaluate our method on Market-1501, Duke-MTMC and CuHK03 datasets. We compare our method with several state-of-the-art methods on the Re-ID task, including Bag of Trick [5], ABDNet [6], Auto-ReID+ [8], TransReID [7], PGFL-KD [9], AML [10], PFD [23], and SOLIDER [3]. For clothes-changing setting, we evaluate our method on LTCC-Standard, LTCC-CC and VC-Clothes datasets. We compare our method with other state-of-the-art methods (i.e., LTCC [2], FSAM [11], CAL [12], GI-ReID [13], AIM [14], and FIRe [15]).
@@ 4.2 Performance Comparisons
We report the performance of our proposed PGS with three different versions: PGS-1, PGS-2, and PGS-3 which include one to three projectors in PHP for transferring posture knowledge. The results for the clothes-consistent problem are shown in Table 1, while the clothes-changing problem results are shown in Table 2.
| Methods | Market-1501 | Duke-MTMC | CuHK03 | |||
| mAP | R1 | mAP | R1 | mAP | R1 | |
| Bag of Trick [5] | 94.2 | 95.4 | 89.1 | 90.3 | 56.6 | 58.8 |
| ABDNet [6] | 88.3 | 95.6 | 78.6 | 89.0 | ||
| Auto-ReID+ [8] | 88.2 | 95.8 | 80.1 | 90.1 | 74.2 | 78.1 |
| TransReID [7] | 89.0 | 95.1 | 82.2 | 90.7 | ||
| PGFL-KD [9] | 87.2 | 95.3 | 79.5 | 89.6 | ||
| AML [10] | 89.5 | 95.7 | 81.7 | 91.1 | 82.3 | 85.6 |
| PFD [23] | 89.7 | 95.5 | 82.2 | 90.6 | ||
| SOLIDER [3] | 95.3 | 96.6 | 86.1 | 89.4 | 71.6 | 67.4 |
| PGS (Our) | 95.4 | 96.9 | 91.4 | 92.6 | 89.7 | 87.9 |
Clothes-consistent: From Table 1, it is evident that our method leads to an improvement of in the mAP metric and in the R1 metric on the Market-1501 dataset when compared to the second-best method, SOLIDER [3]. Regarding the Duke-MTMC dataset, our method achieves a better mAP metric by and an improvement of in the R1 metric compared to the second-best method, Bag of Tricks [5]. In the CuHK03 dataset, our method surpasses the second-best method, AML [10], by a margin of in the mAP metric and in the R1 metric. These results clearly demonstrate that our method is competitive with state-of-the-art approaches across all three datasets, underscoring the effectiveness of leveraging pose information for distinguishing individuals. This outcome is significant in showcasing the potential of our approach for future studies.
@@
| FIRe [15] | 39.9 | 75.9 | 19.1 | 44.6 | ||
| PGS (Ours) | 43.0 | 77.5 | 26.7 | 49.1 | 84.6 | 92.5 |
Clothes-changing: From Table 2, on the LTCC-Standard dataset, our model demonstrates an improvement of over in mAP and in the R1 metric compared to the second-best method, AIM [14]. In the only clothes-changing testing scenario (i.e., LTCC-CC dataset), our method surpasses the second-best (i.e., FIRe) with a improvement in mAP and a substantial margin of in the R1 metric. Regarding the VC-Clothes dataset, our method achieves a significant improvement of in mAP and in R1 metric compared to the second-best method, FSAM. This improvement demonstrates the ability of our method in tackling the clothes-changing problem by focusing on the pose information rather than the clothes.
4.3 Ablation Studies
To evaluate the effectiveness of the PGS framework, we conducted four ablation studies. The first experiment assesses the effectiveness of pose guidance by the PHP module across multiple scales in the clothes-consistent problem, compared to the baseline human encoder (i.e., SOLIDER [3]). The second experiment verify the effectiveness of the PHP module across multiple scales in the clothes-changing problem. The third experiment explores the impact of the parameter , which is used to control the contribution of the guide loss in the total loss, on knowledge transfer. The fourth experiment involves cross-domain testing to examine the generalization capability of our framework.
Effect of PHP across multiple scales in clothes-consistent problem: Table 3 shows that with our proposed PHP module applying to multiple stages, the results are improved over the baseline, even with a relatively small number of added trainable parameters. Since our framework only uses the human encoder in the inference stage, thus the FLOPs value remains the same compared with the baseline.
Impact of for knowledge transfer: As illustrated in Table 5, results indicate that the model’s performance first improves and then declines when the equals 1.0. This phenomenon is due to the fact that the low limits the ability of the model for learning the pose knowledge, while equals 1.0 can make the model difficult to optimize both triplet loss and guide loss.

### 01 Jun 2024 v3

# PGDS: Pose-Guidance Deep Supervision for Mitigating Clothes-Changing in Person Re-Identification

@@ Abstract
Person Re-Identification (Re-ID) task seeks to enhance the tracking of multiple individuals by surveillance cameras. It supports multimodal tasks, including text-based person retrieval and human matching. One of the most significant challenges faced in Re-ID is clothes-changing, where the same person may appear in different outfits. While previous methods have made notable progress in maintaining clothing data consistency and handling clothing change data, they still rely excessively on clothing information, which can limit performance due to the dynamic nature of human appearances. To mitigate this challenge, we propose the Pose-Guidance Deep Supervision (PGDS), an effective framework for learning pose guidance within the Re-ID task. It consists of three modules: a human encoder, a pose encoder, and a Pose-to-Human Projection module (PHP). Our framework guides the human encoder, i.e., the main re-identification model, with pose information from the pose encoder through multiple layers via the knowledge transfer mechanism from the PHP module, helping the human encoder learn body parts information without increasing computation resources in the inference stage. Through extensive experiments, our method surpasses the performance of current state-of-the-art methods, demonstrating its robustness and effectiveness for real-world applications. Our code is available at https://github.com/huyquoctrinh/PGDS.
@@ 1 Introduction
Person Re-Identification (Re-ID) involves tracking and distinguishing individuals across various cameras and viewpoints. This technique finds application in surveillance camera systems, enhancing security measures by enabling the search, tracking, and identifying unidentified individuals within specific areas or towns, thereby enhancing security measures.
Despite notable advancements, several challenges persist in the Re-ID task, e.g. occlusion, weather conditions, and other external factors. In such problems, clothing-change presents itself as one of those paramount challenges, first highlighted by Xue et al. [31] and Wu et al. [30]. This challenge becomes particularly crucial in real-world scenarios where individuals always change their clothes.
In recent years, researchers in the Re-ID domain have increasingly focused on mitigating the clothing-change problem. Various methods, such as VC-Clothes [27], LTCC [25], FSAM [11], CAL [7], and GI-ReID [13], have been introduced to address this issue. The primary difficulty in tackling this challenge arises from deep learning models relying heavily on clothing appearance to make predictions rather than considering biometric traits (e.g., face, hair, pose, gait). A growing body of literature, such as PGFL-KD [33], Pose-transfer [18], Pose-guided re-ID [15], and PSEAMA [14] have demonstrated the effectiveness of pose information in tackling the clothes-chaning problem. However, these works often require complex modules to integrate body parts information into the Re-ID model, leading to an increase in computational cost, training time, and inference time.
As a result, we introduce Pose-Guidance Deep Supervision (PGDS), a simple and innovative framework for the clothes-changing Re-ID task. Figure 1 indicates an example query result of our framework under the clothing-change scenario. In essence, our method builds upon three main modules: human encoder, pose encoder, and Pose-to-Human Projection module (PHP). The human encoder, i.e., the main Re-ID model, efficiently adapts SOLIDER [5], a self-supervised learning framework for human-centric tasks, to extract the general human representations from the input images. At the same time, the pose encoder is taken from the pose estimation model OpenPose [3] to generate meaningful global features for biometric identification, which includes the pose, body, foot, hand, and facial. Our PHP, which includes several projectors, operates between two encoders to transfer the pose information extracted from the pose encoder to the human encoder across various scales. In practice, we fine-tune the pre-trained human encoder and train our projectors while freezing the pose encoder since we only need the general human pose information from the pose encoder. By incorporating pose information into the multiple layers, the Re-ID model can effectively concentrate on the body parts information, which contain unique information crucial for individual identification. Furthermore, integrating the frozen pre-trained pose estimation model also avoids increasing the computational cost in the inference stage, thus benefiting real-world applications. In summary, this paper’s main contributions are as follows:
@@
We introduce a Re-ID framework, termed PGDS, incorporating a frozen pre-trained pose estimation model to guide the human representation model across different scales.
We conduct comprehensive ablation studies on multiple datasets to assess the effectiveness of our proposed PGDS framework.
The content of this paper is organized as follows. In Section 2, we briefly review existing methods for Clothes-changing Re-Identification problem and similar works that explore the utilization of pose information for Person Re-Identification. Then, we introduce our proposed method in Section 3. Section 4 presents the experiment setup, performance comparison, and ablation study. Finally, we conclude our work and suggest problems for future work in Section 5.
@@ 2 Related Work
Clothes-changing Re-Identification: Clothes-changing poses one of the most significant challenges in the Person Re-Identification task. The introduction of the LTCC dataset [25] marked a significant development in addressing the clothes-changing problem. Subsequent research has yielded various methods to tackle this challenge. FSAM [11] aims to obtain coarse ID masks with structure-related details, incorporating ID-relevant information for discriminative structural feature extraction. CAL [7] proposes a novel multi-positive-class classification loss to formulate multi-class adversarial learning. GI-ReID [13] integrates clothes classification and a casual inference model to mitigate bias in clothing information. AIM [32] defines a causality-based auto-intervention model to mitigate clothing bias for robust Clothes-changing Person Re-identification. FIR [28] introduces fine-grained feature mining and a fine-grained attribute recomposition module to enhance the learning of robust features. Although previous works achieved impressive results on the clothes-changing Re-ID, it is worth noting that these methods lack focusing on the crucial parts of the body, which is the salient information to distinguish a specific person. In contrast, our work focus on prioritizing body parts information by efficiently leverage the pose information from frozen pre-trained pose estimation model to guide the main Re-ID model.
Pose Guide for Clothes-changing Re-Identification: Pose is crucial information utilized to guide Re-Identification models. Several methods have been proposed based on this approach. Among the most notable ones are ABDNet [4], PGR [16], Gated Fusion [1], and Pitr [2]. These methods follow the concept of integrating modules to fuse features extracted from pose heatmaps to the Re-ID models. More recently, PGFL-KD [33], PFD [29], Pose-guided re-ID [15], and PSEAMA [14] have leveraged pose guidance through part matching via global features. These methods demonstrate the effectiveness of pose information in guiding Re-ID models. However, the primary issue noted in these approaches is the substantial computational load imposed by the fusion modules, which are used to incorporate pose information into the main model. Differently from these previous works, PGDS integrates the pre-trained pose estimation model during the training phase, which remains frozen, to guide the main Re-ID model with pose information. This approach prevents an increase in computational expenses during the inference stage.
3 Proposed PGDS
In this section, we introduce our PGDS framework, which consists of three primary modules: the human encoder, the pose encoder, and the pose-to-human projection module (PHP), as illustrated in Figure 2.
@@ 3 Proposed PGDS
The human encoder takes in the image, learns each person’s feature representations and unique human appearance information, and then returns the image embedding. Meanwhile, the pose encoder (i.e., from a frozen pre-trained pose estimation model) is used to extract the human pose information from the input image and packs this information into a pose heatmap. In the training phase, the PHP encourages the early layers of the human encoder to attain the human representation with additional pose information through a knowledge distillation mechanism. When inferring, only feature representations from the human encoder are used to accomplish the re-identification task. In other words, we leverage a frozen encoder pre-trained for pose estimation as a teacher to compel the student human encoder to retain pose information into the image embedding. We detail each component in the following subsections to provide more in-depth information.
@@ 3.3 Pose-to-Human Projection Module (PHP)
It is important to note that the human encoder outputs feature vectors that are more informative and semantically richer compared to the pose encoder, which only contains human pose information. From this understanding, it becomes apparent that directly guiding the features from the human encoder with pose features might result in the loss of other crucial information embedded in the human features. Therefore, we design an indirect transfer knowledge mechanism called a projector. This mechanism enables the human encoder to learn pose information without compromising other essential information. Intuitively, the projector acts as a non-linear, lightweight extractor. It unveils human-specific information within the higher-dimensional image representation of the human encoder and maps it to the lower-dimensional pose heatmap space of the pose encoder. Consequently, the resulting features at each stage, denoted as , are incorporated into the guide loss to align them with the pose feature obtained from the pose encoder.
@@ 3.4 Objective Functions
where is the triplet loss [10] to learn person-specific representations, and is the guide loss to transfer the pose knowledge from pose encoder to human encoder. The hyperparameter , which we set to 0.8, controls the contribution and effect of the guide loss on the whole system. See section 5 for the ablation study on the values.
@@ 4.1 Experimental Setup
Datasets: To assess the effectiveness of the proposed PGDS, we conduct evaluations on five datasets that previous methods have widely used. These datasets cover a wide range of scenarios and closely resemble real-world conditions. Market-1501 [34] contains 1,501 identities observed from 6 camera viewpoints, including 12,936 training images of 751 identities, 19,732 gallery images, and 2,228 queries. Duke-MTMC [26] comprises 34,183 images of 1,404 identities from eight cameras. It contains 16,522 training images, 17,661 gallery images, and 2,228 queries. CuHK03 [17] contains 14,097 pictures of 1,467 identities. Additionally, the dataset offers 20 sets for training and testing, where 100 identities are reserved for testing in each set while the remaining identities are used for training. LTCC [25] contains 17,119 person images of 152 identities. The training set in this data includes 9,576 images with 77 identities (46 clothes-changing IDs and 31 clothes-consistent IDs). The testing set includes 493 query images and 7,050 test images with 75 identities (45 clothes-changing IDs + 30 clothes-consistent IDs). This dataset includes two types of testing: Standard (test cases including clothes-changing and clothes-consistent), and CC (test cases include only clothes-changing IDs). VC-Clothes [27] comprises synthesized human images. It encompasses 19,060 images of 512 unique identities in four scenes. In the training phase, 9449 images of 256 identities are employed, leaving the remaining 9611 images of 256 identities for the testing phase.
@@
Implementation Detail: During the training and testing phase, all images are resized to dimensions of . Our training and testing for the clothes-consistent follow the setting of Bag of Trick [22], while in the clothes-changing testing, we follow the setup of LTCC [25]. The Pytorch framework is used, and all experiments are done on the Tesla A100 40GB GPU. The batch size for the training and testing is 64; we also use the learning rate scheduler to support our training experiments with a base learning rate of 8e-4, and the temperature value equals 2. We choose the AdamW[21] as our optimizer, and the best weight has been gotten after 250 epochs. For data augmentation, we employ random horizontal flip and random erasing.
@@
Baselines: Based on two kinds of Re-ID problems, we select different baselines in two scenarios as follows. For clothes-changing setting, we evaluate our method on LTCC-Standard, LTCC-CC and VC-Clothes datasets. We compare our method with other state-of-the-art methods (i.e., LTCC [25], FSAM [11], CAL [7], GI-ReID [13], AIM [32], and FIRe2 [28]). For clothes-consistent setting, we evaluate our method on Market-1501, Duke-MTMC and CuHK03 datasets. We compare our method with several state-of-the-art methods on the Re-ID task, including Bag of Trick [22], ABDNet [4], Auto-ReID+ [6], TransReID [8], PGFL-KD [33], AML [12], PFD [29], and SOLIDER [5]. To further assess the robustness of our model, we conduct the cross-domain setting in two scenarios: the Duke to Market () and the Market to Duke (). We compare our method against several existing methods, e.g., SSL [23], ATNet [19], UDAP [24], and SOLIDER [5].
@@ 4.2 Performance Comparisons
The results of the clothes-changing setting, the clothes-consistent setting, and the cross-domain setting are shown in Table 1, Table 2, and the Table 3, respectively.
@@
| FIRe2 [28] | 39.9 | 75.9 | 19.1 | 44.6 | ||
| PGDS (Ours) | 43.0 | 77.5 | 26.7 | 49.1 | 84.6 | 92.5 |
Clothes-changing: From Table 1, it is evident that on the LTCC-Standard dataset, our model showcases an improvement of over in mAP and in the R1 metric compared to the second-best method, AIM [32]. In the sole clothes-changing testing scenario (LTCC-CC dataset), our method outperforms the second-best (FIRe2) with a improvement in mAP and a substantial margin of in the R1 metric. As for the VC-Clothes dataset, our method achieves a notable improvement of in mAP and in the R1 metric compared to the second-best method, FSAM. These results underscore the efficacy of our method in addressing the clothes-changing problem by prioritizing other body parts over clothing information.
| Methods | Market-1501 | Duke-MTMC | CuHK03 | |||
| mAP | R1 | mAP | R1 | mAP | R1 | |
| Bag of Trick [22] | 94.2 | 95.4 | 89.1 | 90.3 | 56.6 | 58.8 |
| ABDNet [4] | 88.3 | 95.6 | 78.6 | 89.0 | ||
| Auto-ReID+ [6] | 88.2 | 95.8 | 80.1 | 90.1 | 74.2 | 78.1 |
| PGFL-KD [33] | 87.2 | 95.3 | 79.5 | 89.6 | ||
| AML [12] | 89.5 | 95.7 | 81.7 | 91.1 | 82.3 | 85.6 |
| PFD [29] | 89.7 | 95.5 | 82.2 | 90.6 | ||
| SOLIDER [5] | 95.3 | 96.6 | 86.1 | 89.4 | 71.6 | 67.4 |
| PGDS (Our) | 95.4 | 96.9 | 91.4 | 92.6 | 89.7 | 87.9 |
Clothes-consistent: From Table 2, it is evident that our method leads to an improvement of in the mAP metric and in the R1 metric on the Market-1501 dataset when compared to the second-best method, SOLIDER [5]. Regarding the Duke-MTMC dataset, our method achieves a better mAP metric by and an improvement of in the R1 metric compared to the second-best method, Bag of Tricks [22]. In the CuHK03 dataset, our method surpasses the second-best method, AML [12], by a margin of in the mAP metric and in the R1 metric. These results demonstrate that our method remains competitive with state-of-the-art approaches across all three datasets, emphasizing the efficacy of leveraging pose information for distinguishing individuals. Such outcomes are significant as they highlight the potential robustness of our approach for future studies.
Cross-domain testing: From Table 3, in the setting, our method surpasses the second best method by margins of and in terms of mAP and R1, respectively. While in the scenario, we exceed the performance by and in mAP and R1 compared to the second best method. These results highlight the robustness of our framework in the Re-ID task.
4.3 Ablation Study
To evaluate the effectiveness of the PGDS framework, we conducted four ablation studies. The first experiment assesses the effectiveness of pose guidance by the PHP module across multiple scales in the clothes-consistent problem, compared to the baseline human encoder (i.e., SOLIDER [5]). The second experiment verifies the PHP module’s effectiveness across multiple scales in the clothes-changing problem. The third experiment explores the impact of the parameter , which controls the contribution of the guide loss in the total loss, on knowledge transfer.
Effect of PHP across multiple scales in clothes-consistent problem: Table 4 shows that with our proposed PHP module applying to numerous stages, the results are improved over the baseline, even with a relatively small number of added trainable parameters. Since our framework only uses the human encoder in the inference stage, thus the FLOPs value remains the same compared with the baseline.